OpenStudy (anonymous):

How do I find a third vector in the same plane as the first two vectors but at a angle from one of the vectors

OpenStudy (anonymous):

since they are all coplanar vectors, the third vector can be assumed as xi+yj where i and j are unit vectors in the direction of X, Y axes respectively. now you can use the definition of DOT product and solve for x and y.

OpenStudy (amistre64):

 then rotate the vectors as needed to align with the known vectors X = xcos(\(\theta\))-ysin(\(\theta\)) Y = xsin\(\theta\))+ycos(\(\theta\))